OpenStudy (gabylovesyou):

Which of the following is not an equivalent form of the compound inequality x + 18 /= 5 A number line with a closed circle on -13, shading to the left, and a closed circle on -8, shading to the right. 10 >/= x + 18 >/= 5 -13

OpenStudy (helder_edwin):

u have \[ \large x+18\leq10\quad\text{and}\quad x+18\geq5 \] which put togetehr gives u \[ \large 5\leq x+18\leq10 \]

OpenStudy (helder_edwin):

this rules out the second choice

OpenStudy (cwrw238):

yes - but what about the first choice - is that correct?

OpenStudy (helder_edwin):

if u add -18 u get \[ \large -13\leq x\leq -8 \] this rules out the third choice

OpenStudy (pradipgr817):

correct option is a

OpenStudy (cwrw238):

yes a

OpenStudy (helder_edwin):

the 4th choice is